OpenStudy (anonymous):

How many solutions are there to the following system of equations? 2x + 6y = 4 5x + 15y = 10 A. 0 B. 1 C. infinitely many D. 2

OpenStudy (solomonzelman):

\(\large\color{blue}{ 2x + 6y = 4}\) \(\large\color{blue}{ 5x + 15y = 10}\) divide the first equation by 2, and the second equation by 5

OpenStudy (anonymous):

Like this? 2x + 6y divided into 2?

OpenStudy (anonymous):

Will the first one be 1x + 3y = 3 ?

OpenStudy (solomonzelman):

\(\large\color{blue}{ 2x ~~+~~ 6y ~~= ~~4}\) \(\large\color{blue}{ \Downarrow}\) \(\large\color{blue}{ \Downarrow}\) \(\large\color{blue}{ \Downarrow}\) \(\large\color{blue}{ \div2}\) \(\large\color{blue}{ \div2}\) \(\large\color{blue}{ \div2}\) \(\large\color{purple}{ x ~~+~~ 3y ~~= ~~2}\)

OpenStudy (solomonzelman):

do the same division, but by 5, for the second equation.

OpenStudy (anonymous):

Ah-ha! :) Okay, thanks! Let me try to do the second one.

OpenStudy (solomonzelman):

yes, go ahead:)

OpenStudy (anonymous):

@SolomonZelman The second equation is: 2.5x + 7.5y = 5

OpenStudy (solomonzelman):

you are dividing by 5, not by 2

OpenStudy (anonymous):

Right! I forgot. :( The answer is: 1x + 3y = 2

OpenStudy (texaschic101):

if the equations are the same, then there is infinitely many solutions

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

@SolomonZelman What do we do now?

OpenStudy (anonymous):

Oh, I had not noticed that they had the same equations!

OpenStudy (solomonzelman):

\(\large\color{purple}{x+3y=2}\), this is what both of your equations came to be equal to, right?

OpenStudy (anonymous):

Right.

OpenStudy (solomonzelman):

Oh, you get it, i c....

OpenStudy (solomonzelman):

\(\large\color{purple}{x+3y=2}\) \(\large\color{purple}{^{^{\Huge-}}~~x+3y=2}\) \(\large\color{purple}{^{\text{______________}}}\) \(\large\color{blue}{0=0}\)

OpenStudy (solomonzelman):

how many solutions are there?

OpenStudy (solomonzelman):

oh 0=0 is true for any x value, correct?

OpenStudy (anonymous):

There are no solutions.

OpenStudy (anonymous):

I think.

OpenStudy (solomonzelman):

your system, after your divisions that we performed, is: \(\large\color{purple}{x+3y=2}\) \(\large\color{purple}{x+3y=2}\) \(\large\color{purple}{x=2-3y}\) \(\large\color{purple}{x=2-3y}\) \(\large\color{purple}{x=x}\)

OpenStudy (solomonzelman):

x=x has infinitely many solutions.

OpenStudy (solomonzelman):

When ever you have 2 equations that are multiples of each other (and here we do so too, as we have showed), that means infinity of solutions.

OpenStudy (anonymous):

Oh, I think that I understand a little bit better now. :)

OpenStudy (solomonzelman):



OpenStudy (solomonzelman):

two lines on top of each other, they will intersect infinite times

OpenStudy (anonymous):

And so our answer is C. infinitely many.

OpenStudy (solomonzelman):

yup

OpenStudy (anonymous):

Well, thank you for your help and time. I have three more questions, so if I need help can I tag you to the question(s)? @SolomonZelman